OpenStudy (anonymous):

How many pounds are there in 2 and 3 over 4 tons? (1 ton = 2000 pounds) 4075 5000 5500 5575

OpenStudy (anonymous):

StudyGurl14 help

OpenStudy (anonymous):

@StudyGurl14

OpenStudy (studygurl14):

\(\large 2\frac{3}{4}\) Like that?

OpenStudy (anonymous):

yes

OpenStudy (studygurl14):

Do this: \(\large 2\frac{3}{4}\times2000\rightarrow\frac{11}{4}(2000)\rightarrow ???\) Can you do the rest?

OpenStudy (anonymous):

nvm i already know this one sorry i put the wrong one oops

OpenStudy (studygurl14):

Oh, okay. :)